Add the following rational numbers:
Question. "5 by 36 and -7 by 12"
And the correct answer is showing in my book is "-1 by 48" please let me know how is it possible.
Thanks!

Answer by Alan3354(69443)

Question. "5 by 36 and -7 by 12"
-----
Do you mean 5/36 and -7/12 ?
--> 5/36 - 21/36 = -16/36
= -4/9